A manufacturer claims that 10% of women using the "pill" suffer from side effects. The Federal Drug Administration (FDA) believes that the manufacturer's claim is too low and decides to test the manufacturer's claim at α = 5 %. A random sample of 900 women who use the manufacturer's pill shows that 103 suffer side effects. The p-value of this test is:
Answer #1

Solution :

Given that ,

n = 900

x = 103

The null and alternative hypothesis is

H0 : p = 0.10

Ha : p < 0.10

This is the left tailed test .

\hat p = x / n = 103 / 900 = 0.1144

P0 = 10% = 0.10

1 - P0 = 1 - 0.10 = 0.90

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.1144 - 0.10 / [\sqrt0.10 (1 - 0.10) / 900 ]

= 1.444

The test statistic = 1.444

The P-value of this test is : 0.9257